# Disentangled Representation with Cross Experts Covariance Loss for Multi-Domain Recommendation

### v1

@@ 5.4. Ablation Study (RQ3)
Table 6. gAUC metric along with different numbers of embeddings (#Emb) on ME-MMoE-PG and Crocodile. * indicates that the performance difference against ME-MMoE-PG is significant at p-value¡0.05.
#Emb 2 3 4 5 6 7 8 9 10
ME-MMoE-PG 0.66138 0.66157 0.66206 0.66232 0.66203 0.66196 0.66286 0.66279 0.66285
Crocodile 0.66198* 0.66284* 0.66324* 0.66373* 0.6637* 0.66394* 0.66395* 0.66432* 0.66397*
Lift (%) 0.06% 0.13% 0.12% 0.14% 0.17% 0.20% 0.11% 0.15% 0.11%
5.5. Hyper-Parameters and Embedding Sets (RQ4)
@@
In this section, we investigate the sensitivity of the model’s performance to hyperparameters and the number of embeddings. For hyperparameters, the primary parameter is the weight of CovLoss in Eq. 7, i.e., $\alpha$. Taking Kuairand1k as an example, we find that as long as $\alpha$ is not too small (below $2\times 10^{-5}$), Crocodile’s AUC and gAUC remain ¿0.78590 and ¿0.66174, thus maintaining its status as the second-best method.
@@
Additionally, we analyzed the impact of model capacity on performance by adjusting the number of embedding sets. At the minimum number of MEs (i.e., 2-sets), Crocodile achieves a gAUC of 0.66198 and an AUC of 0.78595, which still surpasses other MTL and MDL methods, even when the MTL(ME) methods apply for 5-sets of embeddings. Moreover, as the number of embeddings increases, the performance gains of Crocodile over ME-MMoE-PG are robust ($+0.06\%\sim+0.20\%$), as illustrated in Tab. 6.
5.6. Computational Complexity
@@
The Multi-Embedding (ME) paradigm has already been widely adopted in industry (Zhaok et al., 2021; Ning et al., 2023b; Su et al., 2023; Guo et al., 2023b; Pan et al., 2024). Compared to these ME architectures, our proposed Crocodile introduces the CovLoss upon it, which only influences the training process while not the inference at all. Its complexity is not only small and negligible when compared with the ME-based backbone, but it can be further reduced by sampling.
@@
6. Conclusion
Multi-domain learning is vital in enhancing personalized recommendations. Existing MTL methods fail to learn both dimensional-robust representations and diverse user interests. We propose Crocodile, which employs a Covariance Loss (CovLoss) and Prior informed Element-wise Gating (PEG) based on the multi-embedding (ME) paradigm. Comprehensive experimental evaluations validate the effectiveness of Crocodile.

### v2

@@ 4.5. Ablation Study (RQ3)
4.6. Setting Sensitivity Analysis (RQ4)
@@
In this section, we investigate the sensitivity of the model’s performance to hyper-parameters, for example, the weight of CovLoss $\alpha$ and the number of embeddings. At Kuairand1k, we find that as long as $\alpha$ is not too small (below $2\times 10^{-5}$), Crocodile’s AUC and gAUC remain ¿0.78590 and ¿0.66174, thus maintaining its status as the second-best method.
Figure 8. gAUC metric of Crocodile varying different numbers of embeddings (#Emb) compared with Base, i.e., ME-MMoE with PG, in Kuairand1k. The solid line represents the lift ratio.
@@
Additionally, we analyzed the impact of model capacity on performance by adjusting the number of embedding sets. At the minimum number of MEs (i.e., 2-sets), Crocodile achieves a gAUC of 0.66198 and an AUC of 0.78595, which still surpasses other MTL and MDL methods, even when the other MTL(ME) methods employ 5 embeddings sets. Moreover, as the number of embeddings increases, the performance gains of Crocodile over ME-MMoE with PG are robust ($+0.06\%\sim+0.20\%$). The detailed results are shown in Fig. 8.
4.7. Online Deployment (RQ5)
We conduct an empirical evaluation of Crocodile on Tencent’s online advertising platform. The underlying model architecture utilizes Heterogeneous Experts with Multi-Embedding framework (Pan et al., 2024). Specifically, our approach involves training multiple feature interaction experts, such as GwPFM (Pan et al., 2024) (a variant of FFM (Juan et al., 2016) and FwFM (Pan et al., 2018)), IPNN (Qu et al., 2016), DCN V2 (Wang et al., 2021), or FlatDNN, to capture diverse feature interactions for sparse ID features. Additionally, we learn multiple embedding tables for all features (Guo et al., 2023b; Su et al., 2023; Pan et al., 2018), with each table corresponding to one or several experts. We handle sequence features by TIN (Zhou et al., 2024) to capture the semantic-temporal correlation, numeric features by the multiple Numeral Systems Encoding (Chen et al., 2022; Pan et al., 2024), and embedding features from external models via Similarity Encoding Embedding (Pan et al., 2024).
Specifically, we train a multi-domain model over two main scenarios, with two embedding tables and experts. The output of these two experts is then fed into the primary scenario.
We conduct online A/B testing on March 2024. The baseline model employs a common Embedding & Interaction architecture. During 5% A/B test, we observe a 0.11% AUC lift on several main tasks such as convertible click prediction, and a significant 0.72% CTR lift and 0.73% GMV lift.
False Tolerence
During online training, if the training samples from the auxiliary domains become unavailable or have some issues, we simply roll back the model to the latest valid checkpoint and keep it trained and updated based on only the main domain until the auxiliary one is recovered. If the main domain’s data pipeline has any problems, we roll back the model and keep it fixed.
5. Related Work
5.1. Multi-domain Recommendation
Multi-domain learning (MDL) is a promising topic that has garnered attention from large-scale industrial recommendations, such as Alibaba (Sheng et al., 2021; Zhang et al., 2022; Tian et al., 2023), Meituan (Chang et al., 2023) and Kuaishou (Zhou et al., 2023). Many MDL methods (Sheng et al., 2021; Zhou et al., 2023; Zhang et al., 2022; Li et al., 2023a; Chang et al., 2023; Tian et al., 2023; Wang et al., 2024) have made great efforts to model complex multi-domain interests.
Current methods often start from the Shared-bottom (Caruana, 1997) structure to explore capturing both domain-shared and domain-specific interests. STAR (Sheng et al., 2021) introduces a learnable matrix and dot product operator before outputting the prediction value. HiNet (Zhou et al., 2023) and HAMUR (Li et al., 2023a) allocate domain-shared and domain-specific networks. Furthermore, the PEPNet (Chang et al., 2023) method introduces EPNet, which performs dot product operations between domain-related and domain-shared embeddings, while Maria (Tian et al., 2023) designs feature scaling, refinement, and correlation modeling to capture cross-domain interests at the feature-level. They show the trend of integrating domain information at the more bottom level, i.e., from Tower to Embedding.
Recently, STEM (Su et al., 2023) proposed a multi-embedding (ME) structure in multi-task learning (MTL) and step forward in allocating domain-personalized structures. It achieved better results compared to both single and multi-embedding versions of MMoE (Ma et al., 2018b) and PLE (Tang et al., 2020) in MTL experiments. However, STEM fails in the multi-domain recommendation. We will discuss the reasons in Sec. 2.
Besides, some other paradigms are applied for multi-domain recommendation. For example, AdaSparse (Yang et al., 2022) and ADL (Li et al., 2023b) applied adaptive learning mechanism. M2M (Zhang et al., 2022) utilizes meta-learning techniques. EDDA (Ning et al., 2023a), COAST (Zhao et al., 2023) and UniCDR(Cao et al., 2023) provided graph learning-based solutions.
5.2. Disentangled Representation Learning
Disentangled representation learning helps preserve the distinctiveness of embeddings. Loss-based methods such as OLE (Lezama et al., 2018), dot product (Wu et al., 2021), and cosine similarity (Liu et al., 2021) push representations of different domains towards orthogonality and residing in corresponding subspace. Distance correlation (dCorr) (Liu et al., 2022a) encourages lower dCorr between different factors. ETL (Chen et al., 2023) models the joint distribution of user behaviors across domains with equivalent transformation. VAE-based methods (Ma et al., 2019; Wang et al., 2022; Liu et al., 2022b) promote higher KL divergence between dimensions but often assume a normal prior distribution, which is challenging in the context of sparse interactions (Rendle, 2010), especially with imbalanced sample volumes in multi-domain recommendation. GAN-based models (Chen et al., 2016; Yu and Welch, 2021; Khrulkov et al., 2021) enhance discrimination among representations but are computationally expensive and unsuitable for online recommendation.
Recently, some research has provided systematic and unified insights. VICReg (Bardes et al., 2021; Shwartz-Ziv et al., 2024) uses covariance and variance as loss components to promote independence across dimensions while preventing insufficient learning. Methods like (Chan et al., 2022; Yu et al., 2020) propose a maximal coding rate reduction principle on mixed distributions to maximize the coding rate of the entire representation while minimizing it within each cluster.
5.3. Dimensional Collapse
Dimensional collapse (Chi et al., 2022; Guo et al., 2023b) occurs when all representations span a lower-dimensional space, reflecting the insufficient learning of embeddings. Chi et al. (2022) found that the SMoE structure (Fedus et al., 2022; Lepikhin et al., 2020) may cause dimensional collapse due to gating influences. (Guo et al., 2023b) observed this issue in popular models such as DCN V2 (Wang et al., 2021) and Final MLP (Mao et al., 2023). To measure collapse, they proposed the Information Abundance (IA) (Guo et al., 2023b; Pan et al., 2024) metric, defined as the ratio between the sum and the maximum l1-norm value of the target singular values. A small IA indicates severe insignificance of tail dimensions and potential dimensional collapse.
Some methods (Hua et al., 2021) bridge the gap between representation disentanglement and dimensional collapse prevention. They also claim that feature decorrelation helps alleviate dimensional collapse by reducing strong correlations between axes, promoting more diverse and informative representations.
6. Conclusion
Multi-domain learning is vital in enhancing personalized recommendations. Existing methods fail to learn both dimensional-robust representations and diverse user interests. We propose Crocodile, which employs a Covariance Loss (CovLoss) and Prior informed Element-wise Gating (PEG) based on the multi-embedding (ME) paradigm. Both offline and online experiments validate the effectiveness of Crocodile.
Appendix A Computational Complexity
@@
The Multi-Embedding (ME) paradigm has already been widely adopted in industry (Zhaok et al., 2021; Ning et al., 2023b; Su et al., 2023; Guo et al., 2023b; Pan et al., 2024). Compared to these ME architectures, our proposed Crocodile introduces the CovLoss upon it, which only influences the training process while not the inference at all. Its complexity is not only small and negligible based on the ME-based backbone, but it can be further reduced by only computing the loss based on downsampled samples.